OpenStudy (anonymous):

sin(2arcsin(-1/4)) solve and show steps.

OpenStudy (anonymous):

let x=arcsin(-1/4)=>sin(x)=-1/4...calculate cos(x)... sin(2arcsin(-1/4))=sin(2x)=2sin(x)cos(x) just plug in

OpenStudy (anonymous):

Thanks!